OpenStudy (anonymous):

Use the rational root theorum and the factor theorum to help solve the following equation X^3 + X^2 -25x - 25 = 0

OpenStudy (nenadmatematika):

x^2(x+1)-25(x+1)=(x+1)(x^2-25)=(x+1)(x-5)(x+5)